OpenStudy (shaik0124):

The number of calls coming into a call center on Tuesday are 5/11 more than 1/4 the number that came in on Monday. The total number of calls for the two days together was 270. How many calls came in on Tuesday? a) 50 b) 61 c) 20 d) 72 e) 18

OpenStudy (agl202):

72!

OpenStudy (shaik0124):

how

OpenStudy (shaik0124):

my Approach is let the number of calls on monday be x let the number of calls on tuesday be y then calls on tuesday as per question is y= 5/11 +(x)*1/4 total calls for two days is x+y=270

OpenStudy (shaik0124):

@Agl

OpenStudy (shaik0124):

@ganeshie8

OpenStudy (shaik0124):

@Ashy98

OpenStudy (shaik0124):

any idea?@Ashy98

OpenStudy (shaik0124):

@ashhhrodriguez @eliassaab @FateRocks8D @GracieFerrell @Haseeb96 @iGreen. @k_lynn @Love_Ranaa @MakaylaTracy @Night-Watcher @ParthKohli @Quan99 @rehvi @sammixboo @Teddyiswatshecallsme @undeadknight26 @VectorPrime @waterineyes @x3_x3 @YOLO_KING @zpupster

OpenStudy (yolo_king):

.-.

OpenStudy (shaik0124):

??

OpenStudy (anonymous):

You are good..

OpenStudy (anonymous):

Can you solve those two equations?

OpenStudy (anonymous):

\[y = \frac{5}{11} + \frac{1}{4} x \\ x + y = 270\]

OpenStudy (shaik0124):

@iGreen.

OpenStudy (shaik0124):

@AnswerMyQuestions

OpenStudy (anonymous):

You did not like my reply??